OpenStudy (anonymous):

Find the coordinates of the x-intercept(s) for f(x) = (x^2+10x+24)/(x^2-36)

OpenStudy (anonymous):

I got x=-5

OpenStudy (anonymous):

sorry I actaully got (-4,0)

OpenStudy (anonymous):

(x+6)(x+4) /(x+6)(x-6) = (x+4)/(x+6) = 0 x+4 = 0 x= -4 yep

OpenStudy (anonymous):

yaya!!! thanks!

OpenStudy (anonymous):

croeso

OpenStudy (anonymous):

so it's (-4,0)

OpenStudy (anonymous):

oh yes the y cordinate is -4+4-4+6 = 0

OpenStudy (anonymous):

(-4,0)

OpenStudy (anonymous):

if ur wondering about the 'croeso' - it means welcome